# Relations Among Different Inequality Measures in Complex Systems: From Kinetic Exchange to Earthquake Models

### v1

@@ Abstract
We present a numerical study of several inequality measures across two kinetic wealth-exchange models with extreme inequality features (namely the Banerjee model, and the Chakraborti or Yard-Sale model) and two earthquake simulating models (namely the Chakrabarti-Stinchcombe two-fractal overlap model and the nonlinear dynamical Burridge-Knopoff model), and a synthetic Pareto distribution. For each model we compute numerically the Lorenz function for the respective models’ wealth, overlap magnitude or avalanche distributions. We then estimate the variations of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices in these models with systematic variations of saving propensity (for the two wealth-exchange models), with systematic variations of generation or block numbers (for the two earthquake simulating models). We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models and the two-fractal overlap model) remain a little above unity (theoretically predicted value) and deviating a little higher by a maximum of $4\%$ near $g=k\simeq 0.86$, which was identified earlier to be the precursor point of criticality in several self-organized critical models ($k=0.80$ corresponds to Pareto’s 80-20 law). In the Burridge-Knopoff model for some instances of time, the value of p/(2k1)p/(2k-1) drops a little below unity. This and some other quantitatively similar behaviors of the inequality indices across socio-economic and geophysical models may provide a coherent and comparative framework for identifying the subtle features in the statistics of such disparate dynamical systems.
@@ 1 Introduction
For the socio-economic domain, early models introduced by Dragulescu and Yakovenko [DY_2000] and Chakraborti and Chakrabarti [CC_2000] demonstrated that simple stochastic exchange rules can generate realistic wealth distributions. Extensions involving fixed or distributed saving propensity [CCM_2004], interaction heterogeneity, and asymmetric exchange have yielded richer phenomenology. The Chakraborti Model [C_Model] or also known as Yard-Sale model [YS], later extended by Boghosian and collaborators [YS1], captures multiplicative dynamics and extreme wealth condensation. The Banerjee Model [B_Model, B_Model1] represents another natural kinetic exchange framework, whose intrinsic trading dynamics lead to steady-state configurations characterized by extreme concentration of wealth among a small subset of agents. In this study, we focus on the Banerjee Model, where inequality is tuned by varying the range of exchange and the saving propensity, and the Chakraborti Model (or Yard-Sale model), where saving propensity acts as the central control parameter.
@@
To broaden the scope beyond socio-economic systems, we also examine inequality associated with Pareto-type distributions [Pareto, Pareto1] by tuning their shape parameter $\alpha$, given the ubiquity of power laws in systems exhibiting self-organized criticality. Since such distributions are central to geophysical phenomena, we analyze inequality patterns in two different earthquake models, the Chakrabarti–Stinchcombe two-fractal overlap model [CS_Model, CS1], where event sizes arise from Cantor-set overlaps, and the nonlinear Burridge–Knopoff spring-block model [langer_1994, BK_Model], where block–spring dynamics generate heterogeneous slip events.
@@
Quantifying inequality across these systems , we therefore investigate three Lorenz curve [lor, lor1] based measures: the Gini index (gg[gini], captures the average deviation from perfect equality; The Pietra (or Hoover) index (pp[pietra, pietra1], measures the maximum imbalance in cumulative wealth and the more recently introduced Kolkata index (kk[k-index, k-index1, k-index2, k-index3], identifies the wealth share of the richest fraction of the population and has found applications in diverse socio-economic contexts. For each system, we systematically vary the relevant control parameters such as the exchange range ($R$) and saving propensity ($\lambda$) in the Banerjee Model, saving propensity ($\lambda$) in the Chakraborti Model, the exponent ($\alpha$) in the synthetic Pareto distribution, the Cantor generation index in Chakrabarti–Stinchcombe two-fractal overlap model, and the time evolution of the Burridge–Knopoff model and compute the corresponding inequality indices. In a recent work [asim_2025] authors studied the ratios p/(2k1)p/(2k-1) and $p/g$ in some financial systems. We further examine the ratios p/(2k1)p/(2k-1), $p/g$ for these kinetic exchange models, the two earthquake simulating models, and the Pareto distribution to investigate the variation of these and other relationships among the inequality measures $g$, $p$ and $k$.
@@ 2 Inequality Measures
Quantifying inequality in any distribution whether describing wealth, event sizes, or released energy (hereafter called wealth) requires statistical measures that capture heterogeneity in a systematic manner. In this study, we numerically examine three statistical measures based on the Lorenz function [lor, lor1], namely the Gini index [gini], the Pietra (or Hoover) index [pietra, pietra1], and the Kolkata index [k-index, k-index2].
@@
Here, $x$ denotes the bottom $x$ fraction of the population, and $L(x)$ gives the corresponding fraction of the wealth they hold. For a perfectly equal distribution, $L(x)=x$, producing a 45-degree line known as the perfect equality line (see in Fig. 1). Deviations of $L(x)$ below this line indicate inequality. If the total wealth goes to a single hand, we get the absolute inequality line as seen in Fig. 1.
@@ 2.3 Kolkata Index
In many empirical systems and theoretical models, $k$ displays near-universal scaling behaviour in relation to $g$ or $p$, which motivates our numerically comparative study across two kinetic wealth exchange models and two earthquake simulating models [sand, manna, soumya, asim_2025].
@@ 3.1 Kinetic Wealth-Exchange Models
Kinetic wealth exchange models given by Dragulescu and Yakovenko [DY_2000] describes wealth evolution through pairwise stochastic exchanges among agents and Chakraborti and Chakrabarti through saving propensities [CC_2000]. Below we describes our considered two kinetic wealth exchange models.
@@ 3.1.1 B-Model with Exchange Range
The Banerjee model [B_Model, B_Model1] (or B-Model in short hereafter) introduces a constrained exchange rule in which each agent interacts only within a selected exchange range $R$ in the wealth-ordered list. At each time step $t$:
@@
where $\epsilon$ is a random number drawn uniformly from $(0,1)$. This model naturally generates strong wealth condensation for small $R$, producing highly unequal steady states—though not to the extreme limit observed in the C-Model (discussed later), where all wealth collapses into a single agent. For each choice of $R$, we compute the steady-state Lorenz geometry and associated inequality indices after ensemble averaging.
@@ 3.1.3 C-Model with Savings
The Chakraborti Model [C_Model] (or C-Model in short hereafter) also widely known as Yard-Sale model [YS] is based on multiplicative dynamics in which the amount at stake depends on the wealth of the poorer agent. With saving propensity $\lambda\in[0,1)$, the effective tradable wealth becomes $(1-\lambda)\min(w_{i},w_{j})$. The exchanged amount is $\Delta w=\epsilon(1-\lambda)\min\{w_{i}(t),w_{j}(t)\}$, where $\epsilon$ is uniformly drawn from $(0,1)$. The update equations follow:
@@ 3.2 P-Model
We investigate inequality in synthetic Pareto-type distribution [Pareto, Pareto1] (or P-Model in short hereafter) characterized by the probability density function,
@@
where $p$ denotes the cumulative fraction of the population, and $L(p)$ is the corresponding fraction of total wealth held by that population.
For each $\alpha$ we generate large synthetic datasets, compute the Lorenz curve, and evaluate the associated inequality indices. This baseline serves as a reference for comparing model-generated distributions with known power-law behavior observed in socio-economic and geophysical systems.
@@ 3.3 Earthquake Simulating Models
We now move to two classes of geophysical models that simulate earthquake like event size statistics. The Chakrabarti-Stinchcombe fractal overlap construction generates multifractal, heavy-tailed spatial measures that produce inequality through time variation of the overlap magnitude, whereas the Burridge–Knopoff model gives rise to inequality in terms of spatially varying spring force. Both models probe different but complementary aspects of earthquake dynamics, namely fault geometry and stress distribution during an earthquake. The Chakrabarti–Stinchcombe model exhibits a power-law scaling in the cumulative distribution of overlap magnitudes in the asymptotic limit of large generation number and large overlaps, arising purely from the construction of the underlying fractlas [pbhattacharya]. In contrast, the Burridge-Knopoff model displays power-law event-size statistics in specific dynamical regimes, where nonlinear friction and collective interactions drive the system toward scale-free behavior. Thus, although both models can produce similar phenomenological power-law statistics, the underlying mechanisms are fundamentally different– geometric hierarchy in the former and dynamical evolution in the latter.
@@ 3.3.1 CS-Model
The Chakrabarti Stinchcombe model [CS_Model, CS1, pbhattacharya] (CS-Model in short hereafter) based on the overlap of fractal fault surfaces (modeled using Cantor sets), focuses on fault geometry and the distribution of overlaps that correspond to stored elastic energy. It is therefore a model of geometric heterogeneity: it explains why slip events follow broad distributions (e.g., power laws) purely from the self-similar structure of the fault interface suitable for Lorenz-curve–based inequality analysis. Earthquake magnitudes empirically follow the Gutenberg–Richter law, $\log_{10}N_{r}(M>M_{0})=a-bM$, with $N_{r}(M>M_{0})$ denoting the number of events above magnitude $M_{0}$, and $b\simeq 1$. The model captures this scale-free behavior by treating a fault as two identical self-affine fractals, typically $n^{th}$-generation Cantor sets with $2^{n}$ segments of length $3^{-n}$, sliding over each other. A replica is shifted in discrete steps of $3^{-n}$ under periodic boundary conditions, and the overlap magnitude $Y_{n}(t)$ counts coinciding segments at step $t$, generating a discrete time series $\{Y_{n}(t)\}$. The allowed values of $Y_{n}$ are $2^{\,n-k}$ ($k=0,1,\dots,n$), occurring with exact frequencies
@@
which normalize to the binomial probabilities $\Pr(2^{\,n-k})=\binom{n}{k}\left(\tfrac{1}{3}\right)^{n-k}\left(\tfrac{2}{3}\right)^{k}$; the most probable overlap scales as $\hat{Y}_{n}\sim(2^{n})^{1/3}$. CS-Model captures the static, structural statistics of fault roughness providing insight into how geometric irregularity shape the statistics of earthquakes.
@@ 3.3.2 BK-Model
In contrast, the Burridge–Knopoff Model [langer_1994, BK_Model] (BK-Model in short hereafter) is a dynamical and mechanical model of earthquake generation, where blocks connected by springs and subject to friction evolve over time. It is used to study stick–slip dynamics and stress redistribution. BK-Model captures the temporal dynamics of rupture propagation and seismicity. BK-Model provide insight into how dynamical frictional processes shape the statistics of earthquakes. This framework captures fault dynamics through a chain of $N$ blocks of mass $m$, each connected to its neighbors by springs of stiffness $k_{c}$ and to a slowly driven plate by springs of stiffness $k_{p}$. As the plate moves with velocity $v$, stress accumulates until blocks slip, producing earthquake-like events. Each block follows Newton’s equation of motion,
@@
Using the stick–slip friction law—with parameters $\alpha$ (friction weakening rate) and $\sigma$ (friction drop), the model can reproduce Gutenberg–Richter–like magnitude statistics, exhibit chaotic dynamics despite being deterministic, and transition between periodic, chaotic, and scale-free regimes depending on parameters. Although more sophisticated rate-and-state (R&S) friction laws exist, we use the stick–slip rule which is computationally simpler.
@@ 4 Simulation Framework
For the CS-Model, the fractal representation of a fault surface is constructed using the classical base-3 Cantor set to model the geometrical overlap dynamics between two self-similar rough interfaces. For a given generation $n$, the frequency of each overlap magnitude is evaluated using Eqn. (10), with $n$ ranging from 4 to 200. The inequality measures $g$, $k$, and $p$ are computed from the corresponding Lorenz curves, where the overlap magnitude is plotted along the abscissa and the cumulative frequency along the ordinate. These measures provide a compact quantitative characterization of the degree of heterogeneity in the spatial overlap and enable us to track the evolution of inequality as the fractal resolution is progressively increased.
For the BK model, we perform numerical simulations to investigate the emergence of collective slip dynamics for a given set of physical parameters. The simulations are carried out under prescribed frictional and elastic conditions, specified by the friction coefficient $\mu$, the spring constants $k_{p}$, and the number of blocks $N$. At each time step, the instantaneous force acting on each block is computed using Eqns. (11) and (12), and the corresponding Lorenz curve is constructed by plotting the force magnitude along the x-axis, and the cumulative fraction of blocks along the y-axis. The degree of heterogeneity in the force distribution is quantified using inequality indices, which serve as compact descriptors of the non-uniformity in the stress landscape and provide a time-resolved measure of the buildup and release of inequality within the block ensemble.

### v2

@@ Abstract
We present a numerical study of several inequality measures across two kinetic wealth-exchange models with extreme inequality features (namely the Banerjee model, and the Chakraborti or Yard-Sale model) and two earthquake simulating models (namely the Chakrabarti-Stinchcombe two-fractal overlap model and the nonlinear dynamical Burridge-Knopoff model). For each model we compute numerically the Lorenz function for the respective models’ wealth, overlap magnitude or avalanche distributions. We then estimate the variations of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices in these models with systematic variations of saving propensity (for the two wealth-exchange models), with systematic variations of generation or block numbers (for the two earthquake simulating models). We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models and the two-fractal overlap model) remain a little above unity (theoretically predicted value) and deviating a little higher by a maximum of $4\%$ near $g=k\simeq 0.86$, which was identified earlier to be the precursor point of criticality in several self-organized critical models ($k=0.80$ corresponds to Pareto’s 80-20 law). The Burridge–Knopoff model shows larger deviations from the $k$-$g$ relation, but does not access the high-inequality regime where an intersection occurs. This and some other quantitatively similar behaviors of the inequality indices across socio-economic and geophysical models may provide a coherent and comparative framework for identifying the subtle features in the statistics of such disparate dynamical systems.
@@
Inequality-how unevenly outcomes are distributed, shows up not only in economics, but also in many driven physical systems where typically fewer larger events (avalanches) tend to dominate the major scaling behavior. In this work, we use a common quantitative language to compare ‘extreme inequality’ across two kinetic wealth-exchange models and two earthquake-inspired models. From the simulated distributions (wealth, fractal-overlap magnitudes, and slip avalanches), we construct the Lorenz curve and extract three standard inequality measures: the Gini index ($g$), the Pietra index ($p$), and the Kolkata index ($k$). By systematically varying key control parameters in each model (saving propensity in the economic models; system size or generation/block number in the geophysical models), we show that a specific composite ratio, p/(2k1)p/(2k-1), stays close to its theoretical benchmark—typically slightly above unity and deviating by at most $\sim 8\%$, generally near a previously identified precursor or threshold regime around $g=k\approx 0.86$. Overall, these cross-model regularities suggest that inequality indices can serve as sensitive, comparable diagnostics for recognizing subtle changes in the statistics of disparate complex systems, from markets to fault dynamics.
@@ 1 Introduction
For the socio-economic domain, early models introduced by Dragulescu and Yakovenko [16] and Chakraborti and Chakrabarti [11] demonstrated that simple stochastic exchange rules can generate realistic wealth distributions. Extensions involving fixed or distributed saving propensity [13], interaction heterogeneity, and asymmetric exchange have yielded richer phenomenology. The Chakraborti Model [12] or also known as Yard-Sale model [8], later extended by Boghosian and collaborators [7], captures multiplicative dynamics and extreme wealth condensation. The Banerjee Model [5, 18] represents another natural kinetic exchange framework, whose intrinsic trading dynamics lead to steady-state configurations characterized by extreme concentration of wealth among a small subset of agents. In this study, we focus on the Banerjee Model, where inequality is tuned by varying the range of exchange and the saving propensity, and the Chakraborti Model (or Yard-Sale model), where saving propensity acts as the central control parameter.
@@
To broaden the scope beyond socio-economic systems, we also examine inequality associated with Pareto-type distributions [28, 26] by tuning their shape parameter $\alpha$, given the ubiquity of power laws in systems exhibiting self-organized criticality. Since such distributions are central to geophysical phenomena, we analyze inequality patterns in two different earthquake models, the Chakrabarti Stinchcombe two-fractal overlap model [10, 23], where event sizes arise from Cantor-set overlaps, and the nonlinear Burridge Knopoff spring-block model [9, 27], where block–spring dynamics generate heterogeneous slip events.
@@
Quantifying inequality across these systems , we therefore investigate three Lorenz curve [24, 1] based measures: the Gini index (gg[21], captures the average deviation from perfect equality; The Pietra (or Hoover) index (pp[29, 17], measures the maximum imbalance in cumulative wealth and the more recently introduced Kolkata index (kk[20, 14, 4, 3], identifies the wealth share of the richest fraction of the population and has found applications in diverse socio-economic contexts. For each system, we systematically vary the relevant control parameters such as the exchange range ($R$) and saving propensity ($\lambda$) in the Banerjee Model, saving propensity ($\lambda$) in the Chakraborti Model, the exponent ($\alpha$) in the synthetic Pareto distribution, the Cantor generation index in Chakrabarti Stinchcombe two-fractal overlap model, and the time evolution of the Burridge Knopoff model and compute the corresponding inequality indices. In a recent work [19] authors studied the ratios p/(2k1)p/(2k-1) and $p/g$ in some financial systems. We further examine the ratios p/(2k1)p/(2k-1), $p/g$ for these kinetic exchange models, the two earthquake simulating models, and the Pareto distribution to investigate the variation of these and other relationships among the inequality measures $g$, $p$ and $k$.
@@ 2 Inequality Measures
Quantifying inequality in any distribution whether describing wealth, event sizes, or released energy (hereafter called wealth) requires statistical measures that capture heterogeneity in a systematic manner. In this study, we numerically examine three statistical measures based on the Lorenz function [24, 1], namely the Gini index [21], the Pietra (or Hoover) index [29, 17], and the Kolkata index [20, 4].
@@
Here, $x$ denotes the bottom $x$ fraction of the population, and $L(x)$ gives the corresponding cumulative fraction of the wealth they hold. For a perfectly equal distribution, $L(x)=x$, producing a 45-degree line known as the perfect equality line (see in Fig. 1). Deviations of $L(x)$ below this line indicate inequality. If the total wealth goes to a single hand, we get the absolute inequality line as seen in Fig. 1.
@@ 2.3 Kolkata Index
In many empirical systems and theoretical models, $k$ displays near-universal scaling behaviour in relation to $g$ or $p$, which motivates our numerically comparative study across two kinetic wealth exchange models and two earthquake simulating models [2, 25, 15, 19].
@@ 3.1 Kinetic Wealth-Exchange Models
Kinetic wealth exchange models given by Dragulescu and Yakovenko [16] describes wealth evolution through pairwise stochastic exchanges among agents and Chakraborti and Chakrabarti through saving propensities [11]. Below we describes our considered two kinetic wealth exchange models.
@@ 3.1.1 B-Model with Exchange Range
The Banerjee model [5, 18] (or B-Model in short hereafter) introduces a constrained exchange rule in which each agent interacts only within a selected exchange range $R$ in the wealth-ordered list. At each time step $t$:
@@
where $\epsilon$ is a random number drawn uniformly from $(0,1)$. This model naturally generates strong wealth condensation for small $R$, producing highly unequal steady states, though not to the extreme limit observed in the C-Model (discussed later), where all wealth collapses into a single agent. For each choice of $R$, we compute the steady-state Lorenz geometry and associated inequality indices after ensemble averaging.
@@ 3.1.3 C-Model with Savings
The Chakraborti Model [12] (or C-Model in short hereafter) also widely known as Yard-Sale model [8] is based on multiplicative dynamics in which the amount at stake depends on the wealth of the poorer agent. With saving propensity $\lambda\in[0,1)$, the effective tradable wealth becomes $(1-\lambda)\min(w_{i},w_{j})$. The exchanged amount is $\Delta w=\epsilon(1-\lambda)\min\{w_{i}(t),w_{j}(t)\}$, where $\epsilon$ is uniformly drawn from $(0,1)$. The update equations follow:
@@ 3.2 P-Model
We investigate inequality in synthetic Pareto-type distribution [28, 26] (or P-Model in short hereafter) characterized by the probability density function,
@@
where $p$ denotes the fraction of the population, and $L(p)$ is the corresponding cumulative fraction of total wealth. For each $\alpha$ we generate large synthetic datasets, compute the Lorenz curve, and evaluate the associated inequality indices. This baseline serves as a reference for comparing model-generated distributions with known power-law behavior observed in socio-economic and geophysical systems.
@@ 3.3 Earthquake Simulating Models
We now move to two classes of geophysical models that simulate earthquake like event size statistics. The Chakrabarti Stinchcombe fractal overlap construction generates multifractal, heavy-tailed spatial measures that produce inequality through time variation of the overlap magnitude, whereas the Burridge Knopoff model gives rise to inequality in terms of spatially varying spring force. Both models probe different but complementary aspects of earthquake dynamics, namely fault geometry and stress distribution during an earthquake. The Chakrabarti Stinchcombe model exhibits a power-law scaling in the cumulative distribution of overlap magnitudes in the asymptotic limit of large generation number and large overlaps, arising purely from the construction of the underlying fractals [6]. In contrast, the Burridge Knopoff model displays power-law event-size statistics in specific dynamical regimes, where nonlinear friction and collective interactions drive the system toward scale-free behavior. Thus, although both models can produce similar phenomenological power-law statistics, the underlying mechanisms are fundamentally different, geometric hierarchy in the former and dynamical evolution in the latter.
@@ 3.3.1 CS-Model
The Chakrabarti Stinchcombe model [10, 23, 6] (CS-Model in short hereafter) based on the overlap of fractal fault surfaces (modeled using Cantor sets), focuses on fault geometry and the distribution of overlaps that correspond to stored elastic energy. It is therefore a model of geometric heterogeneity: it explains why slip events follow broad distributions (e.g., power laws) purely from the self-similar structure of the fault interface suitable for Lorenz-curve–based inequality analysis. Earthquake magnitudes empirically follow the Gutenberg–Richter law,
$\log_{10}N_{r}(M>M_{0})=a-bM,$
with $N_{r}(M>M_{0})$ denoting the number of events above magnitude $M_{0}$, and $b\simeq 1$. The model captures this scale-free behavior by treating a fault as two identical self-affine fractals, typically $n^{th}$-generation Cantor sets with $2^{n}$ segments of length $3^{-n}$, sliding over each other. A replica is shifted in discrete steps of $3^{-n}$ under periodic boundary conditions, and the overlap magnitude $Y_{n}(t)$ counts coinciding segments at step $t$, generating a discrete time series $\{Y_{n}(t)\}$. The allowed values of $Y_{n}$ are $2^{\,n-k}$ ($k=0,1,\dots,n$), occurring with exact frequencies
@@
which normalize to the binomial probabilities
$\Pr(2^{\,n-k})=\binom{n}{k}\left(\tfrac{1}{3}\right)^{n-k}\left(\tfrac{2}{3}\right)^{k};$
the most probable overlap scales as $\hat{Y}_{n}\sim(2^{n})^{1/3}$. CS-Model captures the static, structural statistics of fault roughness providing insight into how geometric irregularity shape the statistics of earthquakes.
@@ 3.3.2 BK-Model
In contrast, the Burridge Knopoff Model [9, 27] (BK-Model in short hereafter) is a dynamical and mechanical model of earthquake generation, where blocks connected by springs and subject to friction evolve over time. It is used to study stick–slip dynamics and stress redistribution. BK-Model captures the temporal dynamics of rupture propagation and seismicity. BK-Model provide insight into how dynamical frictional processes shape the statistics of earthquakes. This framework captures fault dynamics through a chain of $N$ blocks of mass $m$, each connected to its neighbors by springs of stiffness $k_{c}$ and to a slowly driven plate by springs of stiffness $k_{p}$. As the plate moves with velocity $v$, stress accumulates until blocks slip, producing earthquake-like events. Each block follows Newton’s equation of motion,
@@
Using the stick–slip friction law—with parameters $\delta$ (friction weakening rate) and $\sigma$ (friction drop), the model can reproduce Gutenberg–Richter–like magnitude statistics, exhibit chaotic dynamics despite being deterministic, and transition between periodic, chaotic, and scale-free regimes depending on parameters. Although more sophisticated rate-and-state (R&S) friction laws exist, we use the stick–slip rule which is computationally simpler.
@@ 4 Simulation Framework
For the CS-Model, the fractal representation of a fault surface is constructed using the classical base-3 Cantor set to model the geometrical overlap dynamics between two self-similar rough interfaces. For a given generation $n$, the frequency of each overlap magnitude is evaluated using Eqn. (10), with $n$ ranging from 4 to 200. The inequality measures $g$, $k$, and $p$ are computed from the corresponding Lorenz curves, where the cumulative normalized overlap magnitude is plotted along the ordinate against the cumulative fraction of events along the abscissa. These measures provide a compact quantitative characterization of the degree of heterogeneity in the spatial overlap and enable us to track the evolution of inequality as the fractal resolution is progressively increased.
For the BK-Model, we perform numerical simulations to investigate the emergence of collective slip dynamics for a given set of physical parameters. The simulations are carried out under prescribed frictional and elastic conditions, specified by the friction coefficient $\mu$, the spring constants $k_{p}$, and the number of blocks $N$. At each time step, the instantaneous force acting on each block is computed using Eqns. (11) and (12), and the corresponding Lorenz curve is constructed by plotting the force magnitude along the y-axis, and the cumulative fraction of blocks along the x-axis. The degree of heterogeneity in the force distribution is quantified using inequality indices, which serve as compact descriptors of the non-uniformity in the stress landscape and provide a time-resolved measure of the buildup and release of inequality within the block ensemble.